Session: Sep 2000 (session 3)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Huang Xinyong at Sat Sep 16 05:38:09 METDST 2000) Success

Effelsberg (A. Kraus at Sat Sep 16 18:12:56 METDST 2000) Success

Westerbork (Tony Foley at Mon Sep 18 09:41:13 METDST 2000) Minor failures
OK from 00:00 to 02:30, with 14 telescopes in the array, but lost 2 telescopes (RT1 and RT9) about 02:30 because of a problem caused by lightning strike.

Torun (K.Borkowski at Mon Sep 18 13:45:39 METDST 2000) Success
No known problems.

Onsala (A. Polatidis at Mon Sep 18 23:43:17 METDST 2000) Success

Urumqi (Liu Xiang at Wed Sep 20 10:31:48 METDST 2000) Success

Noto (C.Trigilio at Wed Sep 20 14:12:05 METDST 2000) Success

Jodrell_Bank (Alastair Gunn at Fri Sep 22 09:48:45 METDST 2000) Success
No known problems.

Cambridge (Alastair Gunn at Fri Sep 22 09:49:27 METDST 2000) Minor failures
Problems with tape deck at start of experiment. Set-up for second scan at 00:12 UT. Second scan may be also corrupt, noticed wrong IF input on BBCs. Everything seems fine after second scan.